suppose that permanent income is calculated as the average of income over the past 5 years
1
Expert's answer
2021-09-07T16:32:39-0400

complete question

solution

a)

"YP=\frac{1}{5}(Y_{t}+Y_{t-1}+Y_{t-2}+Y_{t-3}+Y_{t-4})\\\frac{1}{5}\times (20000+20000+20000+20000+20000)\\=\$20000"

b)

"YP=\frac{1}{5}(Y_{t+1}+Y_1+Y_{t-1}+Y_{t-2}+Y_{t-3}+Y_{t-4})\\\frac{1}{5}\times (30000+20000+20000+20000+20000)\\=\$22000"